John McCarthy Computer Science Department Stanford University jmc@cs.stanford.edu

http://www-formal.stanford.edu/jmc/ 1997 March 27

FORMAL REASONING GROUP Usable Concepts of Logical AI

• Contexts as objects

• Elaboration tolerance

• Reasoning about action

• Challenges

Contexts as objects

Glue together independently written knowledge bases, planning systems, etc.

• ist(c, p), value(c, term)

• value(AF db, price(GE771)) = $6 million • enter(AF db) [later leave(AF db)]

• afdb : price(GE771) = $6 million •

outer : value(AF db, price(GE771))
= value(GEdb, price(GE771))
+price(AF -warranty(GE771))

Elaboration tolerance—preliminary • Some formalisms tolerate elaboration bet-

ter.

• Some elaborations are easier.

• Split a concept, e.g. altitude of an airplane • MYCIN—How to add time? More than one

patient? Contexts.

• Drosophila is Missionaries and Cannibals. – an oar on each bank

– small missionary and large cannibal – more missionaries and cannibals – an island

– one missionary is Jesus Christ

Reasoning about action Compound actions (preliminary)

• Conditional actions. Do X if p.

• Do X until p. Bomb the bridge until it is destroyed.

• Do X for all x in A. Wash all the dishes.

• Washing a dish is not a planned action for humans. We don’t think fast enough for that.

What knowledge should an HPKB have about the effects of compound actions?

Challenges Elaboration tolerance of knowledge formalisms

• Add new preconditions for actions.

• Remove preconditions for actions.

• Add effects of actions.

• Make effects conditional.

• Split a concept.

• Add parameters to a function or predicate, e.g. “It depends on how much they know about it.”